OpenStudy (anonymous):

PLEASE HELLLPPP!! I will give medal! Two cards are drawn from a standard deck of cards. Part A: If they are drawn with replacement, what is the probability that both cards are 2s? Part B: If they are drawn without replacement, what is the probability that the first card is a club and the second card is a spade? Part C: Which of the two scenarios in Part A or Part B represents dependent events? Explain.

OpenStudy (anonymous):

What about it do you not understand?

OpenStudy (anonymous):

I kind of need everything explained. Maybe go step by step @legendarysadist

OpenStudy (anonymous):

Very few people on here will go through a step-by-step explanation of everything. A base knowledge of the subject is first required before you ask a certain question. If it is probability you are confused with, go review it before asking questions. If it is the wording of the question, what about it confuses you?

OpenStudy (anonymous):

The probability is what confuses me, how exactly do I find that?

OpenStudy (anonymous):

Probability is amount you want/amount total. If there are 10 cards, and 5 are red, 5 are blue, what is your chance of getting a red card?

OpenStudy (anonymous):

50%?

OpenStudy (anonymous):

Yes. So if there are 13 cards, and 1 of them is a 2, what is your chance of getting a 2?

OpenStudy (anonymous):

6%?

OpenStudy (anonymous):

You can just keep it in fraction form, so 1/13 (and it would be 8% anyways)

OpenStudy (anonymous):

Now let's say you return the 2 back to the deck, what is your chance of getting it again?

OpenStudy (anonymous):

Ok so if you have 2 cards out of 52, with replacement, does that mean you divide 2/52?

OpenStudy (anonymous):

If you are taking a card twice, that means you multiply the probabilities of getting it both times.

OpenStudy (anonymous):

I don't want to get you confused, so let's start at the beginning

OpenStudy (anonymous):

In a standard deck, there are 4 2's and 52 cards total. So what would that ratio be simplified?

OpenStudy (anonymous):

It would be 2?

OpenStudy (anonymous):

4/52 simplifies to what?

OpenStudy (anonymous):

7% or 0.07

OpenStudy (anonymous):

AGAIN KEEP IT AS A FRACTION. WE ARE NOT DIVIDING.

OpenStudy (anonymous):

Okay sorry 1/13

OpenStudy (anonymous):

As a reminder, we are going to stay in fractions always, unless the question asks for the answer as a decimal or percentage. Fraction should be the default. But let's continue

OpenStudy (anonymous):

okay got it :)

OpenStudy (anonymous):

So the chance of getting a 2 is 1/13. Now if we put the card back into the deck, will we have the same chance of getting the 2 or a different one?

OpenStudy (anonymous):

We will have the same chance.

OpenStudy (anonymous):

Good. So if we want to get the chance of having both happen, what we do is multiply the probability. So 1/13 * 1/13 = ?

OpenStudy (anonymous):

so 1/169

OpenStudy (anonymous):

So this is if they are drawn with replacement?

OpenStudy (anonymous):

Correct. That is our answer for the first problem. Yes, that is right.

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

So what differences do you think there will be in solving it when we do the second problem?

OpenStudy (anonymous):

Since it's without replacement, there will be a change in the probability

OpenStudy (anonymous):

Correct. Anything else?

OpenStudy (anonymous):

I am not sure, this is the one that really confused me.

OpenStudy (anonymous):

Well do you know how many spades there are in a deck? What about clubs?

OpenStudy (anonymous):

there are 13 I think

OpenStudy (anonymous):

That is correct. So what would the probability be of getting a spade out of all of the cards in the deck?

OpenStudy (anonymous):

so would we do 13/52?

OpenStudy (anonymous):

That would find us the probability of getting a spade, yes.

OpenStudy (anonymous):

1/4 would be the probability

OpenStudy (anonymous):

Correct. Now we took 1 card out of the deck, so how many cards in the deck are there now? (Since we aren't replacing the card)

OpenStudy (anonymous):

13/51

OpenStudy (anonymous):

That doesn't simplify any further either.

OpenStudy (anonymous):

Good. So now we will multiply the probabilities together.

OpenStudy (anonymous):

13/204 would be the final answer?

OpenStudy (anonymous):

Yes, that would be it

OpenStudy (anonymous):

Oh wow that made it so much easier. I have a few more, do you have time to help?

OpenStudy (anonymous):

Sure, I can do a couple more

OpenStudy (anonymous):

A manufacturer produces soda cans and a quality control worker randomly selects two cans from the assembly line for testing. Past statistics show that 10% of the cans are defective. What is the probability that the two selected cans are defective if the quality control worker selects the two cans from a batch of 60 cans? P(Both defective) = 6/25 P(Both defective) = 1/118 P(Both defective) = 3/250 P(Both defective) = 9/625

OpenStudy (anonymous):

Any ideas on how you would do this?

OpenStudy (anonymous):

We can start with 2/60?

OpenStudy (anonymous):

10% of all the cans are defective. So for 60 cans, how many are defective?

OpenStudy (anonymous):

6 cans are defective

OpenStudy (anonymous):

So for the first can, what is the chance of it being a defective can?

OpenStudy (anonymous):

6/60 = 1/10

OpenStudy (anonymous):

Good. Now for the second can, we are now missing 1 can out of the defective cans, and one can out of the total cans. So what would the new probability be for the second can?

OpenStudy (anonymous):

5/59? I am not sure about that one

OpenStudy (anonymous):

Then we multiply

OpenStudy (anonymous):

Well you got it, so no worries. Now multiply the two probabilities together to get your answer :)

OpenStudy (anonymous):

1/118 is the answer! Thanks. Here is the next one! Consider two events such that P(A) = 1/4, P(B) = 2/3, and P(A ∩ B) = 1/5. Are events A and B independent events? Yes, they are independent because P(A) ⋅ P(B) = P(A ∩ B) No, they are dependent because P(A) ⋅ P(B) = P(A ∩ B) Yes, they are independent because P(A) ⋅ P(B) ≠ P(A ∩ B) No, they are dependent because P(A) ⋅ P(B) ≠ P(A ∩ B)

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

Okay so if they are independent, that means that they are with replacement, right?

OpenStudy (anonymous):

Yes, that is pretty much what it means.

OpenStudy (anonymous):

Well the first two choices are out because 1/4 * 2/3 does not equal 1/5

OpenStudy (anonymous):

@legendarysadist

OpenStudy (anonymous):

Yes, you are on the right track

OpenStudy (anonymous):

So how do I know if it's dependent or independent?

OpenStudy (anonymous):

Well I could explain this for a while but simply, if you cannot multiply them together then that means something must be changed in them, so they must be dependent of each other.

OpenStudy (anonymous):

Oh I got it! You have no idea how much you are helping me! Thank you thank you thank you So I just have one more! You still have time?

OpenStudy (anonymous):

Yeah I can do one more

OpenStudy (anonymous):

At an amusement park, the probability that a child eats a hot dog and drinks a soda pop is 0.48. The probability that a child eats a hot dog is 0.59, and the probability that a child drinks soda pop is 0.88. What is the probability (rounded to the nearest hundredth) that a child drinks soda pop given that the child has already eaten a hot dog? 0.55 0.81 0.28 0.42

OpenStudy (anonymous):

Just multiply the probabilities together.

OpenStudy (anonymous):

I think it's the third choice. 0.28

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

I think I misread

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Ok, so the child eating a hot dog is for certain, so we will not deal with it. The two variables we have left are that he will drink soda, and the chance of him doing both. So we will multiply them together.

OpenStudy (anonymous):

0.88 * 0.59?

OpenStudy (anonymous):

.88 * .48

OpenStudy (anonymous):

No it's 0.48*0.88

OpenStudy (anonymous):

and the answer is 0.42

OpenStudy (anonymous):

Yes, that is correct. The way they ask this one is a bit confusing tbh

OpenStudy (anonymous):

Yeah tell me about it! Thanks for all your help!!!!!

OpenStudy (anonymous):

Haha no problem (: